OpenStudy (anonymous):

WHY ME? Find the inverse of f(x) = 2x^2 - 3. Explain each step and verify that it is an inverse.

OpenStudy (anonymous):

\[x=2y ^{2}-3\]

OpenStudy (arbershabani97):

just replace x with y so x=2y^2-3 x+3=2y^2 (x+3)/2=y^2 \[y=\sqrt{(x+3)/2}\]